OpenStudy (anonymous):

translate the statement into symbols: the number of hours it takes to paint a house is inversely proportional to the numbers of people painting

OpenStudy (anonymous):

let t= number of hours and n= number of workers

OpenStudy (mathstudent55):

a and b are directly proportional: a = kb a and b are inversely proportional: a = k/b

OpenStudy (anonymous):

so would it be t=k/n?

OpenStudy (anonymous):

or t=n/k?

OpenStudy (mathstudent55):

You need t (hours) in terms of n (workers), so the equation is: t = k/n, where k is the constant of porportionality

OpenStudy (mathstudent55):

Now let's look at it and reason it out. Think of k as a constant number that does not change. Since n is in the denominator, as n bcomes larger, the fraction on the right becomes smaller. Right?

OpenStudy (mathstudent55):

@aij83j Are you still there?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Did you read my last answer?